OpenStudy (anonymous):

Find all the real-number roots of the equation. Give an exact expression for the root. e^x+e^-x=2

OpenStudy (zarkon):

\[e^x-2+e^{-x}=0\] multiply through by \[e^x\] \[e^{2x}-2e^x+1=0\] let \[u=e^x\] then we have \[u^2-2u-1=0\] solve for u...return back to x

OpenStudy (zarkon):

mistake: should be \[u^2-2u+1=0\]

OpenStudy (anonymous):

why can't i just subtract [e ^{x}\]

OpenStudy (zarkon):

?

OpenStudy (anonymous):

e^x

OpenStudy (anonymous):

i have an idea. solve \[z+\frac{1}{z}=2\] and the replace \[z=e^x\]

OpenStudy (anonymous):

\[z+\frac{1}{z}=2\] \[z^2+1=2z\] \[z^2-2z+1=0\] \[(z-1)^2=0\] \[z=1\] \[e^x=1\] \[x=0\]